# Causal Inference with Missing Exposures and Missing Outcomes

### v2

@@ Abstract
Missing data are ubiquitous in public health research. When estimating causal effects, there are well-established methods to address bias to due censored outcomes. Commonly, causal estimands are defined under hypothetical interventions to “set” the exposure and to prevent censoring. Identification is evaluated with the sequential backdoor criterion and considerations of data support. Then inverse weighting, standardization, and doubly-robust approaches are applied for statistical estimation and inference. We demonstrate how this framework can be extended to settings with missingness on the exposure of interest as well as the variable defining the population of interest (e.g., persons at risk of the outcome). Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda. This study posed several real-world challenges: confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at risk of TB), and missingness on the outcome at follow-up (capturing who acquired TB). We present a series of causal models and identification results to demonstrate the handling of missing exposures and outcomes in prospective studies. We highlight the use of TMLE with Super Learner and the real-world consequences of our approach.
Keywords: Causal Inference, Missing Data, Real-world Evidence, Super Learner, Targeted Minimum Loss-based Estimation, TMLE, Tuberculosis
@@ Introduction
Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products [1, 2, 3, 4, 5, 6, 7, 8]. There is rich history of methods research to address the potential for bias when participants with measured outcomes differ meaningfully from those with missing or censored outcomes (e.g., [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]). The Causal Roadmap provides one such approach [21, 22, 23]. First, we specify our causal question: how would expected outcomes differ if all did versus did not receive the exposure of interest and censoring were prevented. Then, we specify a causal model, such as directed acyclic graph (DAG) or non-parametric structural equation model (NPSEM), to represent the data generating process for observed data: baseline and time-varying confounders, exposures, and outcomes [10]. Third, we intervene on the causal model to generate counterfactual outcomes under hypothetical interventions to set the exposures and ensure outcome measurement. Fourth, we evaluate whether the corresponding causal estimand can be identified (i.e., if sequential exchangeability and positivity hold) [9, 10, 11, 12]. Fifth, we specify the statistical estimand, which is often a complex function of the observed data distribution (i.e., not equal to single regression coefficient). Sixth, we conduct estimation and inference with inverse weighting, standardization, or doubly-robust methods, such as targeted minimum loss-based estimation (TMLE) [11]. Finally, we conduct sensitivity analyses and appropriately interpret the results.
@@
Here, we provide a tutorial on how this framework can be extended to address missingness on the exposure. We also highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest. Suppose, for example, we are interested in studying the incidence of some disease. Our population of interest would be persons who are at risk of the developing the outcome and are, thereby, disease-free at baseline. In this setting, our incidence estimates would be subject to bias if there is differential missingness of outcomes at baseline. Using Counterfactual Strata Effects [24, 5, 25, 26, 27, 28, 29], we provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios.
We note there is a growing interest in the use of missingness graphs to represent studies with missingness on multiple variables and to assess whether causal effects can be identified (termed “recovered”) [30, 31]. In particular, Moreno-Betancur and colleagues introduced complete-data DAGs (c-DAGs), missingness DAGs (m-DAGs), and their link [4]. They also provided a series of “canonical” m-DAGs for causal effects of point-treatment exposures [4, 32]. Holovchak et al. extended this work for the effects of longitudinal exposures [33], while questioning the practical utility of the m-DAG approach. Specifically, they noted, “no general algorithms are available to decide on recoverability, and decisions have to be made on a case-by-case basis” [33]. As an alternative, we demonstrate how an established framework for causal inference can used to define, identify, and estimate causal estimands, such as the average treatment effect, under missingness on exposures and outcomes. In the following, we build-up from simple to complex examples in hope that our structured presentation is relevant a wide range of readers who need to address confounding and multi-source missingness. We illustrate practical relevance with the SEARCH-TB study.
Motivating example
@@
SEARCH was a cluster randomized trial to evaluate a community-based approach to a Universal HIV Test-and-Treat intervention, as compared to the standard-of-care, in rural Kenya and Uganda (2013-2017; NCT01864603) [34]. Following a rapid census, all communities were offered multi-disease testing through community health campaigns with home-based follow-up for non-participants [35]. Through this mechanism, we measured demographic data (e.g., age, sex, education, and mobility), self-reported alcohol use (our primary exposure of interest), and tested for HIV infection. Due to high costs and complex logistics, evaluation of incident tuberculosis (TB) infection was limited to SEARCH-TB, a sub-study in 9 eastern Ugandan communities [36, 37]. This sub-study was intentionally enriched for persons with HIV. Specifically, in each community, we sampled 100 households with at least one adult (15+ years) with HIV and 100 households without an adult with HIV. At baseline of the sub-study, tuberculin skin tests were administered to residents of the sampled households. One year later, follow-up tests were administered to participants who tested negative at baseline. Here, we demonstrate the methods used to evaluate the effect of alcohol use on incident TB infection [38]: confounding, missingness on the exposure of interest, missingness on the baseline outcome (defining who was at risk of TB), and missingness on the final outcome (defining who acquired TB).
Related Causal Problems: Building Complexity
@@
Many studies feature only a subset of the challenges described above. We, thus, provide causal models and identification results for a series of hypothetical studies with increasing complexity in the hope of providing a useful reference for a broad range of real-world studies. For simplicity, we focus on defining and identifying causal parameters under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or causal risk ratio).
Classic point-treatment problem
In Appendix S1, we review the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. This could represent a study of the effect of alcohol use ($A$) on incident TB infection ($Y$) among a representative cohort of persons without TB at baseline. Let $Y^{a}$ be the counterfactual outcome if, possibly contrary-to-fact, the participant had exposure-level $A=a$. The counterfactual mean outcome $\mathbb{E}(Y^{a})$ is identified if there are no unmeasured confounders and sufficient data support: $Y^{a}\perp A\mid L$ and $\mathbb{P}(A=a\mid L)>0$, respectively. The statistical estimand is given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid A=a,L)]$[9]. Even if these assumptions are not reasonable, the G-computation formula is still a well-defined statistical estimand, on which we can focus our estimation efforts (Appendix S1).
@@
Most real-world studies, however, depart from this idealized point-treatment problem. Here, we review causal models and identification when there is missingness on the outcome. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not test for TB at the end of follow-up. Let $\Delta_{Y}$ be an indicator of outcome measurement. If $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as usual. However, if $\Delta_{Y}=0$, their outcome $Y$ is not observed. In Figure 1, we show two possible ways of using causal models to represent such a study. In Panel A, we introduce $Y^{\circ}$ to represent the underlying value of the outcome, and define the observed outcome as the product of its measurement indicator and underlying value: $Y=\Delta_{Y}\times Y^{\circ}$. Thus, $Y$ equals its underlying value $Y^{\circ}$ when it’s measured (i.e., when $\Delta_{Y}=1$) and is zero otherwise (i.e., when $\Delta_{Y}=0$). In Panel B, we omit the underlying outcome and directly represent the causal model in terms of the observed data: $O=(L,A,\Delta_{Y},Y)$. In the latter presentation, the measurement indicator $\Delta_{Y}$ again influences the value of the observed outcome $Y$, because when an outcome is not measured, it is set deterministically to NA. Both approaches lead us to same identification assumptions and statistical estimands. Therefore, to minimize notation and mirror the long-standing literature on censoring (e.g., [16, 17, 11, 18, 19, 20, 12]), we use the latter representation in the remainder of the article.
@@
To define causal effects when the outcome is subject to missingness, we now consider counterfactuals indexed by both the exposure and the outcome measurement indicator. Specifically, let $Y^{*}=Y^{A=a,\Delta_{Y}=1}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were at level $A=a$ and missingness on the outcome prevented. Then our causal target parameter $\mathbb{E}(Y^{*})$ is the counterfactual mean outcome if all participants had exposure $A=a$ and their outcomes measured. To identify this causal parameter and express it as a function of the observed data distribution, we would need the baseline covariates $L$ to be sufficient to control for confounding and differential outcome missingness. Specifically, we need $L$ to capture all the common causes of the exposure and outcome, and, among those with the exposure of interest, all the common causes of the outcome and its measurement. These assumptions can be represented as $Y^{*}\perp A\mid L$ and $Y^{*}\perp\Delta_{Y}\mid A=a,L$, respectively. These conditions are often referred to as the “sequential randomization assumption” or “sequential exchangeability” and can be evaluated graphically through the sequential backdoor criterion [9, 10, 11, 12]. We also need a positive probability of being exposed to level $A=a$ within all possible values of the confounders and, among those with the exposure of interest, a positive probability of outcome measurement within all possible values of the confounders: $\mathbb{P}(A=a\mid L)>0$ a.e. and $\mathbb{P}(\Delta_{Y}\mid A=a,L)>0$ a.e., respectively. If these assumptions hold, we have equivalence between our causal estimand $\mathbb{E}(Y^{*})$ and the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,A=a,L)]$ [9].
@@
Figure 1: Causal models with missingness on the outcome (Panels A-B) or missingness on the exposure (Panel C): LL=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $Y^{\circ}$=underlying outcome, $\Delta_{Y}$=indicator of outcome measurement, YY= observed outcome. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
@@
We now demonstrate how the above approach for missing outcomes can be applied to missing exposures. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not answer questions about their alcohol use. Let $\Delta_{A}$ be an indicator that a participant has their exposure measured. If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual. However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is set deterministically to NA. The DAG and NPSEM for such a study are given in Figure 1C. To define causal effects when the exposure is subject to missingness, we consider counterfactuals indexed by both the exposure and its measurement indicator. Now, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and at level $A=a$.
Then to identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we need analogous conditions as the prior subsection on missing outcomes. Specifically, we need $L$ to be sufficient to control for differential exposure missingness and for confounding (among those with measured exposures): $Y^{*}\perp\Delta_{A}\mid L$ and $Y^{*}\perp A\mid\Delta_{A}=1,L$, respectively. These assumptions can be graphically evaluated with the sequential backdoor criterion. We also need the two analogous assumptions on data support: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e. and $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e. If these four assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in Appendix S2.1.
@@
We now combine our two challenges: missing exposures and missing outcomes. Continuing our running example, suppose that among our cohort of persons at risk of TB, some participants did not answer questions about their alcohol use and, despite best efforts, some participants could not be found at the end of the study for outcome ascertainment. To reflect this data generating process, we introduce new notation to reflect the longitudinal data setting. Let $L_{0}$ be baseline covariates and $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment. Figure 2 provides the causal models for such a study.
@@
To define the causal effect when the exposure and outcome are subject to missingness, we now consider counterfactuals indexed by the exposure and two measurement indicators. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$ denote the counterfactual outcome under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement. To identify the counterfactual mean outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders. Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship. Therefore, we rely on sequential randomization/exchangeability and find a set of covariates that satisfies the backdoor criterion for each “intervention” node given the observed past [9]. As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding. In other words, we need the analogous identification assumptions given in the prior subsection. Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$), the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement as well as a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$ and $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e., respectively. If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in Appendix S2.2 [9, 39, 40].
Missing Exposures and Missing Outcomes at the Start and End of Follow-up
@@
We now consider missingness on the outcome at the start of follow-up. Following our motivating example, we are interested in the effect of alcohol use on incident TB infection, but did not reach 100% of study participants for baseline TB testing. In other words, our longitudinal cohort of participants without TB is subject to selection bias. To reflect this data generating process, we update our notation to have multiple outcome measures. Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at start of follow-up (hereafter “baseline”) and $\Delta_{Y_{1}}$ be an indicator of outcome measurement at the end of follow-up (hereafter “endline”). Let $Y_{0}$ and $Y_{1}$ denote the corresponding outcomes; they are set to NA when not observed. The corresponding causal models can be found in Figure 3. The time-ordering reflects the study protocol and procedures in SEARCH-TB; specifically, alcohol use $A$ was measured before baseline TB status $Y_{0}$. As a result, the exposure $A$ can impact who has prevalent TB at baseline ($Y_{0}=1$) and, thereby, who is at risk of TB at baseline ($Y_{0}=0$).
@@
Figure 3: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome at baseline and endline: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, A=exposure, $\Delta_{Y_{0}}$=indicator of outcome measurement at baseline, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of outcome measurement at endline, and $Y_{1}$= outcome at endline. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
@@
To define the causal target parameter in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$. Additionally, we consider the counterfactual outcome at endline under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at endline among those known to be at risk at baseline. Concretely, only participants who tested TB-negative at baseline ($\Delta_{Y_{0}}=1$ and $Y_{0}=0$) were approached for re-testing at endline. Thereby, this final intervention is a dynamic or personalized one to set $\Delta_{Y_{1}}$ equal to one if $Y_{0}=0$ and to zero otherwise (e.g., [24, 41, 42, 43]). For simplicity, denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
@@
Now we can precisely define the causal parameter in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among those at risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$. Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects”[24, 5, 25, 26, 27, 28], which are defined by Nakato et al. as “causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure” [29]. These effects are different from Principal Strata Effects, which are defined within subgroups of latent classes that are fundamentally not observable [44, 45, 46, 47]. For example, in SEARCH-TB, principal stratification could be applied to define the effect of alcohol use on incident TB infection among the subset of participants who would have always tested regardless of their alcohol use. Instead, our interest is the effect of alcohol use on incident TB among the entire population of persons at risk at baseline.
To identify this effect, we re-express the conditional probability as

### v3

@@ Abstract
Missing data are ubiquitous in public health research. When estimating causal effects, there are well-established methods to address bias to due missing outcomes. Commonly, causal estimands are defined under hypothetical interventions to “set” the exposure and to prevent missingness. We demonstrate how this framework can be extended to missing exposures. We further extend this framework to incorporate missingness on the baseline outcome, which induces missingness on the population of interest. To do so, we highlight the use of Counterfactual Strata Effects: causal estimands where the focus population is subject to missingness and/or impacted by the exposure. Our work is motivated by SEARCH-TB’s investigation of the effect of alcohol consumption on the risk of incident tuberculosis (TB) infection in rural Uganda. This study posed several real-world challenges: confounding, missingness on the exposure (alcohol use), missingness on the baseline outcome (defining who was at-risk of TB and, thus, in the focus population), and missingness on the outcome at follow-up (capturing who acquired TB). We present a series of causal models and identification results to demonstrate the handling of missingness in these settings. We highlight the use of TMLE with Super Learner and the real-world consequences of our approach.
Keywords: Causal Inference, Counterfactual Strata Effects, Missing Data, Real-world Evidence, Super Learner, Targeted Minimum Loss-based Estimation, TMLE, Tuberculosis
@@ Introduction
Missing data affect the integrity of analyses across the spectrum of public health research, including surveillance studies to estimate disease prevalence and randomized trials to establish efficacy of new medical products.[37, 80, 69, 46, 5, 12, 33, 42] There is rich history of methods research to address the potential for bias when participants with measured outcomes differ meaningfully from those with missing or censored outcomes.[61, 52, 76, 28, 31, 67, 73, 59, 70, 68, 77, 84] The Causal Roadmap provides one such approach.[53, 15, 14, 21] The standard implementation of the Roadmap with censored outcomes is as follows. First, we specify our causal question: how would expected outcomes differ if all did versus did not have the exposure of interest and censoring were prevented. Then, we specify a causal model, such as a directed acyclic graph (DAG) or non-parametric structural equation model (NPSEM), to represent the data generating process for the baseline and time-varying confounders, exposures, and outcomes.[52, 76] Third, we intervene on the causal model to generate counterfactual outcomes under hypothetical interventions to set the exposures and ensure outcome measurement. Fourth, we evaluate whether the corresponding causal estimand can be identified (i.e., if sequential exchangeability and positivity hold).[61, 52, 76, 28] Fifth, we specify the statistical estimand, which is often a complex function of the observed data distribution (i.e., not equal to single regression coefficient). Sixth, we conduct estimation and inference with inverse weighting, standardization, or doubly robust methods, such as targeted minimum loss-based estimation (TMLE).[76] Finally, we conduct sensitivity analyses and appropriately interpret the results.
@@
Here, we first discuss how this framework for missing and censored outcomes can be extended to address missingness on the exposure. We then introduce Counterfactual Strata Effects: causal estimands defined for a group which is subject to missingness and/or influenced by the exposure.[7, 54, 2, 25, 5, 4, 50, 40, 55, 22, 47] Specifically, we highlight the consequences of missingness on the baseline outcome when it is crucial to defining the population of interest. Concretely, suppose we are interested in studying the incidence of some disease. Our focus population would be persons who are at-risk of the developing the outcome and are, thereby, disease-free at baseline. In this setting, our incidence estimates would be subject to bias if our analyses did not account for differential missingness of outcomes at baseline. We provide a framework for explicitly defining, identifying, and estimating parameters in such scenarios—overall and, importantly, when the exposure influences the baseline risk of the outcome. We also discuss how our approach avoids “failures” sometimes arising in the target trial emulation framework.[26]
@@
We illustrate practical relevance with SEARCH-TB, which aimed to evaluate the effect of alcohol use on incident tuberculosis (TB) infection.[1] SEARCH-TB was a sub-study of SEARCH, a cluster randomized trial to evaluate a community-based approach to Universal HIV Test-and-Treat in rural Kenya and Uganda (2013-2017; NCT01864603).[25] At baseline in SEARCH, we conducted a rapid census and then population-based measurement of sociodemographic factors (e.g., age, sex, mobility, and alcohol use) and HIV.[11] Due to high costs and complex logistics, evaluation of incident TB infection was limited to eastern Ugandan and intentionally enriched for persons with HIV.[39, 40] Specifically, we over-sampled households with at least one adult with HIV. Then we aimed to administer tuberculin skin tests to residents of the sampled households. One year later, we aimed to administer follow-up tests to participants who tested negative at baseline in the sub-study. Here, we demonstrate the methods used to address several real-world challenges arising in SEARCH-TB: confounding, missingness on the exposure of interest (self-report of any alcohol use), missingness on the baseline outcome (defining who was at-risk of TB), and missingness on the final outcome (defining who acquired TB).
Building to Counterfactual Strata Effects
@@
Many studies feature only a subset of the challenges described above. We provide causal models and identification results for a series of hypothetical studies with increasing complexity. For simplicity, we focus on defining and identifying causal estimands under a single level of the exposure, but our results naturally generalize to causal effects defined in terms of contrasts of counterfactual outcome distributions under two levels of the exposure (i.e., the average treatment effect or causal risk ratio). In Appendix S1, we review the classic “point-treatment” problem, where we have measured confounding by baseline covariates $L$, a binary exposure $A$ occurring at single time-point, and an outcome $Y$ occurring at the study’s close. Appendix S2 provides a brief review of approaches to handle missingness on the baseline confounders $L$.
@@
Consider our motivating study of the effect of alcohol use on incident TB infection. Suppose that we have a representative cohort of persons without TB at baseline, but some participants did not test for TB at the end of follow-up. Let $\Delta_{Y}$ be an indicator of outcome measurement. If $\Delta_{Y}=1$ for a participant, we observe their outcome $Y$ as usual. However, if $\Delta_{Y}=0$, their outcome $Y$ is not observed. Figure 1 provides two possible causal models to represent such a study. In Panel A, we use $\Upsilon$ to represent the underlying value of the outcome, and define the observed outcome as the product of its measurement indicator and underlying value: $Y=\Delta_{Y}\times\Upsilon$. Thus, $Y$ equals its underlying value $\Upsilon$ when it’s measured (i.e., when $\Delta_{Y}=1$) and is zero otherwise (i.e., when $\Delta_{Y}=0$). In Panel B, we omit the underlying outcome and directly represent the causal model in terms of the observed data: $O=(L,A,\Delta_{Y},Y)$. In the latter presentation, the measurement indicator $\Delta_{Y}$ again influences the value of the observed outcome $Y$, because the outcome is missing if $\Delta_{Y}=0$. Both approaches lead us to same identification assumptions and statistical estimands. Therefore, to minimize notation and mirror the long-standing literature on censoring, we use the latter representation in the remainder of the article.[60, 59, 8, 70, 76, 68, 77, 84, 28]
@@
To define causal effects when the outcome is subject to missingness, we consider counterfactuals indexed by both the exposure and the outcome measurement indicator. Specifically, let $Y^{*}=Y^{A=a,\Delta_{Y}=1}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were at level $A=a$ and missingness on the outcome prevented. Then our causal estimand $\mathbb{E}(Y^{*})$ is the expected counterfactual outcome if all participants had exposure $A=a$ and their outcomes measured. To identify this causal estimand and express it as a function of the observed data distribution, we would need the baseline covariates $L$ to be sufficient to control for confounding and differential outcome missingness. Specifically, we need $L$ to capture all the common causes of the exposure and outcome, and, among those with the exposure of interest, all the common causes of the outcome and its measurement. These assumptions can be represented as $Y^{*}\perp A\mid L$ and $Y^{*}\perp\Delta_{Y}\mid A=a,L$, respectively. These conditions are often referred to as the “sequential randomization assumption” or “sequential exchangeability” and can be evaluated graphically through the sequential backdoor criterion.[61, 52, 76, 28] We also need a positive probability of being exposed to level $A=a$ within all possible values of the confounders and, among those with the exposure of interest, a positive probability of outcome measurement within all possible values of the confounders: $\mathbb{P}(A=a\mid L)>0$ a.e. and $\mathbb{P}(\Delta_{Y}\mid A=a,L)>0$ a.e., respectively. If these assumptions hold, we have equivalence between our causal estimand $\mathbb{E}(Y^{*})$ and the statistical estimand given by the G-computation formula: $\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,A=a,L)]$.[61]
@@
Figure 1: Causal models with missingness on the outcome (Panels A-B) or missingness on the exposure (Panel C): LL=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, AA=observed exposure, $\Upsilon$=underlying outcome, $\Delta_{Y}$=indicator of outcome measurement, and YY= observed outcome. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
@@
We now demonstrate how the above approach for missing outcomes can be applied to missing exposures. Continuing our running example, suppose that among our representative cohort of persons without TB at baseline, some participants did not answer questions about their alcohol use. Let $\Delta_{A}$ be an indicator that a participant has their exposure measured. If $\Delta_{A}=1$ for a participant, we observe their exposure $A$ as usual. However, if $\Delta_{A}=0$ for a participant, their exposure $A$ is missing. The DAG and NPSEM for such a study are given in Figure 1C. To define causal effects when the exposure is subject to missingness, we consider counterfactuals indexed by both the exposure and its measurement indicator. Now, let $Y^{*}=Y^{\Delta_{A}=1,A=a}$ be the counterfactual outcome for a given participant if, possibly contrary-to-fact, their exposure were measured and at level $A=a$.
Then to identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L,\Delta_{A},A,Y)$, we need analogous conditions as the prior subsection. Specifically, we need $L$ to be sufficient to control for differential exposure missingness and for confounding (among those with measured exposures): $Y^{*}\perp\Delta_{A}\mid L$ and $Y^{*}\perp A\mid\Delta_{A}=1,L$, respectively. We also need the two analogous assumptions on data support: $\mathbb{P}(\Delta_{A}=1\mid L)>0$ a.e. and $\mathbb{P}(A=a\mid\Delta_{A}=1,L)>0$ a.e. If these four assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ as the statistical estimand $\mathbb{E}[\mathbb{E}(Y\mid A=a,\Delta_{A}=1,L)]$ with proof in Appendix S3.1.
@@
We now combine our two challenges: missing exposures and missing outcomes. Continuing our running example, suppose that among our cohort of persons at-risk of TB, some participants did not answer questions about their alcohol use and some participants did not test for TB at the end of follow-up. To reflect this data generating process, we introduce new notation to reflect the longitudinal setting. Let $L_{0}$ be baseline covariates and $L_{1}$ be additional covariates collected after the exposure but before outcome ascertainment. Figure 2 provides the causal models for such a study.
@@
To define the causal effect when the exposure and outcome are subject to missingness, we now consider counterfactuals indexed by the exposure and two measurement indicators. Specifically, let $Y^{*}=Y^{\Delta_{A}=1,A=a,\Delta_{Y}=1}$ denote the counterfactual outcome under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement. To identify the expected counterfactual outcome $\mathbb{E}(Y^{*})$ and express it as a function of the distribution of the observed data $O=(L_{0},\Delta_{A},A,L_{1},\Delta_{Y},Y)$, we now need to account for the post-baseline covariates $L_{1}$, which act as time-dependent confounders. Specifically, $L_{1}$ are mediators of the exposure-outcome relationship, while “confounding” the measurement-outcome relationship. Therefore, we rely on sequential randomization/exchangeability and find a set of covariates that satisfies the backdoor criterion for each “intervention” node given the observed past.[61] As before, we need that the baseline covariates $L_{0}$ are sufficient to control for missing exposures and for confounding. In other words, we need the analogous identification assumptions given in the prior subsection. Additionally, we need that among participants with measured exposures at the level of interest (i.e., $\Delta_{A}=1$ and $A=a$), the baseline and time-varying covariates ($L_{0},L_{1})$ capture all the common causes of outcomes and their measurement as well as a positive probability of outcome measurement within all possible values of the baseline and time-varying covariates: $Y^{*}\perp\Delta_{Y}\mid L_{1},A=a,\Delta_{A}=1,L_{0}$ and $\mathbb{P}(\Delta_{Y}=1\mid L_{1},A=a,\Delta_{A}=1,L_{0})>0$ a.e., respectively. If these assumptions hold, we can rewrite $\mathbb{E}(Y^{*})$ in terms of the longitudinal G-computation formula: $\mathbb{E}\{\mathbb{E}[\mathbb{E}(Y\mid\Delta_{Y}=1,L_{1},A=a,\Delta_{A}=1,L_{0})\mid A=a,\Delta_{A}=1,L_{0})]\}$, shown in terms of iterated expectations and with proof in Appendix S3.2.[61, 8, 72]
Counterfactual Strata Effects
Following our motivating example, we now introduce missingness on the outcome at baseline, which induces missingness on the focus population. Specifically in SEARCH-TB, we were interested in the effect of alcohol use among all persons at-risk of TB, but we did not reach all for initial TB testing. Therefore, our longitudinal cohort of participants who tested TB-negative at baseline in the sub-study was not representative of all persons at-risk of TB. Furthermore, alcohol use likely influenced baseline TB status.[38, 58, 56] Under the target trial framework, SEARCH-TB would result in “Emulation failure 2: time zero is set at eligibility but after strategy assignment”.[26] However, we now demonstrate how following the Causal Roadmap facilitates clear definition of the causal estimand with Counterfactual Strata Effects, while avoiding bias due to missing data during estimation and inference.
@@
To do so, we update our notation to have multiple outcome measures. Let $\Delta_{Y_{0}}$ be an indicator of outcome measurement at start of follow-up (hereafter “baseline”) and $\Delta_{Y_{1}}$ be an indicator of outcome measurement at the end of follow-up (hereafter “endline”). Let $Y_{0}$ and $Y_{1}$ denote the corresponding outcomes. The corresponding causal models can be found in Figure 3. The time-ordering reflects the study protocol and procedures in SEARCH-TB; specifically, alcohol use $A$ was measured before baseline TB status $Y_{0}$.[39, 40, 1] As a result, the exposure $A$ can impact who has prevalent TB at baseline ($Y_{0}=1$) and, thereby, who is at-risk of TB at baseline ($Y_{0}=0$).
@@
Figure 3: Causal graph and corresponding non-parametric structural equation model with missingness on the exposure and the outcome at baseline and at endline: $L_{0}$=baseline covariates, $\Delta_{A}$=indicator of exposure measurement, A=exposure, $\Delta_{Y_{0}}$=indicator of outcome measurement at baseline, $Y_{0}$=baseline outcome, $L_{1}$=time-dependent covariates, $\Delta_{Y_{1}}$=indicator of outcome measurement at endline, and $Y_{1}$= outcome at endline. For ease of presentation, the models are shown without dependence between the unmeasured variables, which are omitted on the graph.
@@
To define the causal estimand in this setting, we first consider the counterfactual outcome at baseline under hypothetical interventions to ensure exposure measurement, “set” the exposure level to $A=a$, and ensure outcome measurement at baseline: $Y_{0}^{*}=Y_{0}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1}$. Additionally, we consider the counterfactual outcome at endline under the prior interventions as well as a hypothetical intervention to ensure outcome measurement at endline among those known to be at-risk at baseline. Concretely, only participants who tested TB-negative at baseline ($\Delta_{Y_{0}}=1$ and $Y_{0}=0$) were approached for re-testing at endline. Thereby, this final intervention is a dynamic or personalized.[7, 29, 74, 62] We set $\Delta_{Y_{1}}$ equal to one if $Y_{0}=0$ and to zero otherwise. For simplicity, denote the resulting counterfactual outcome as $Y_{1}^{*}=Y_{1}^{\Delta_{A}=1,A=a,\Delta_{Y_{0}}=1,\Delta_{Y_{1}}=1}$.
@@
Now we can precisely define the causal estimand in terms of the following conditional probability, which captures the counterfactual incidence of the outcome among all at-risk at baseline: $\mathbb{P}(Y_{1}^{*}=1\mid Y_{0}^{*}=0)$. Due to conditioning on a counterfactual variable, such parameters are sometimes called “Counterfactual Strata Effects”, which are defined by Nakato et al. as “causal estimands where the outcome is only relevant for a group whose membership is subject to missingness and/or impacted by the exposure”.[47] These estimands have previously been specified to evaluate intervention effects in cluster randomized trials.[7, 54, 2, 25, 5, 4, 50, 40, 55, 22, 47] To our knowledge, this is their first use in observational studies and with additional missingness on the exposure.
To facilitate identification, we re-express the conditional probability as